OpenStudy (anonymous):

Riley finds out that Cooper’s Storage and Shipping Company is working with a local business to package some office supplies. Some of the supplies are packed inside a cube-shaped box with side lengths of 4 1/2 in. These boxes are then packed into a shipping box with dimensions of 18 in. × 9 in. × 4 1/2 in. How many boxes of office supplies can be packed into the larger box for shipping?

OpenStudy (anonymous):

@KamiBug :)

OpenStudy (kamibug):

First find the volume of a cube-shaped box then find the volume of the box they go into. Remember the sides of a cube are all the same length. 4.5*4.5*4.5 = 21.125 18*9*4.5 = 729 Now divide the volume of the larger box by the volume of the smaller box to see how many can go in it. :) 729 / 21.125 = ?

OpenStudy (anonymous):

34?

OpenStudy (kamibug):

\(\Huge\cal\color{indigo}{You~got~it!~:)}\)

OpenStudy (anonymous):

thanks:)